If you have available several 2.0-uF capacitors, each capable of withstanding 200V without breakdown, how would you assemble a combination having an equivalent capacitance of (a) 0.40 uF or of (b) 1.2 uF, each capable of withstanding 1000 V?


1
Expert's answer
2021-11-01T17:13:42-0400

Now use 5 capacitor in series combination

"\frac{1}{c}=\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}=\frac{5}{2}"

"C=\frac{2}{5}=0.4\mu F"

Now above this series line make these three line in parallel combination

"C_{net}=c_1+c_2+c_3"

"c_{net}=0.4+0.4+0.4=1.2\mu F"